Consider a population consisting of the values (1, 3, 8), n= 2 with replacement.


  1. mean of sampling distribution of means
  2. variance of the sampling distribution of means
  3. standard deviation of the sampling distribution
1
Expert's answer
2021-05-07T09:33:13-0400

Solution:

Possible sample of size 2 with replacement"=\{(1,3)(3,8)(1,8),(1,1),(3,3),(8,8)\}"

Sample means for these samples"=\{2,5.5,4.5,1,3,8\}"

(1): Mean of sample means"=(2+5.5+4.5+1+3+8)\/6=4"

Population mean"=(1+3+8)\/3=4"

Population variance",\sigma^2=\dfrac{(1-4)^2+(3-4)^2+(8-4)^2}{3}=\dfrac{26}3"

(2): The variance of the sampling distribution of means",\sigma_x^2=\dfrac{\sigma^2(N-n)}{n(N-1)}"

"=\dfrac{\dfrac{26}3(3-2)}{2(3-1)}=\dfrac{13}6"

(3): The standard deviation of the sampling distribution",\sigma_x=\sqrt{\dfrac{13}6}"